OpenStudy (anonymous):

f'' of : f(x)= (1-3x)^3 (2x-3) Please :)

OpenStudy (nincompoop):

so what's the problem here?

OpenStudy (nincompoop):

I can teach you the power rule, but first I want you to show me how you would do this using the derivative definition

OpenStudy (anonymous):

I can't arrive to the answerrrrr

OpenStudy (nincompoop):

tic toc tic toc tic toc

OpenStudy (nincompoop):

hey, nin! what's up, man? do you need help in math?

OpenStudy (nincompoop):

I guess you don't need help anymore @aml13 ?

OpenStudy (anonymous):

no yesss

OpenStudy (nincompoop):

I think the reason you're having a little trouble arriving at the correct answer is because you didn't simplify your function first

OpenStudy (nincompoop):

but I mean it should be still the same utilizing the power rule \[f'(x)= (nx^{n-1})\]

OpenStudy (e.mccormick):

Looks like a mistake in the cain rule.

OpenStudy (e.mccormick):

chain

OpenStudy (nincompoop):

so let us say we have f(x) = 2x^4 it's derivative would be f'(x) = 4(2x^3) = 8x^3

OpenStudy (anonymous):

yes but that's what I did

OpenStudy (e.mccormick):

Hmmm...

OpenStudy (anonymous):

i'm doing f''

OpenStudy (e.mccormick):

Kk. Then... how did you get that? It looks like the mitake is there.

OpenStudy (e.mccormick):

Take each side of the + seperately as you do that part.

OpenStudy (shamil98):

yeah what mick said, mistake with product rule

OpenStudy (anonymous):

its not u'v+uv' ??

OpenStudy (nincompoop):

product (-3+2 x) (d/dx((1-3 x)^3))+(1-3 x)^3 (d/dx(-3+2 x))

OpenStudy (shamil98):

f(x)= (1-3x)^3 (2x-3) get the first derivative.. f'(x) = 2(1-3x)^2(2x-3) + (1-3x)^3(2) f''(x) = 4(1-3x)(2x-3)(2-6x)^3) + 2(1-3x)^2(2x-3)(3(2-6x)^2)

OpenStudy (e.mccormick):

It is u'v+uv', but that is not what you did.

OpenStudy (shamil98):

its messy so i might've made a mistake..

OpenStudy (nincompoop):

I had to write it down laughing out loud

OpenStudy (shamil98):

I think i did it right lol

OpenStudy (anonymous):

haha that's good :)

OpenStudy (anonymous):

Do u want to know the real answer?

OpenStudy (anonymous):

f''= 18(1-3x)(12x-11)

OpenStudy (shamil98):

oh oops i used the product rule again

OpenStudy (shamil98):

no wonder LOL

OpenStudy (nincompoop):

>.<

OpenStudy (nincompoop):

still need help ?

OpenStudy (shamil98):

f'(x) = 2(1-3x)^2(2x-3) + (1-3x)^3(2) f''(x) = 4(1-3x)(2x-3) + 6(1-3x)^2

OpenStudy (e.mccormick):

And Nin, was the other way you were hinting at the multiply it all out method?

OpenStudy (nincompoop):

why are you using 2 sham?

OpenStudy (nincompoop):

laughing out loud mic, that was dirty and loads of work

OpenStudy (e.mccormick):

Hehe. Well... it does "simplify" the problem.

OpenStudy (shamil98):

the derivative of (2x-3) is just 2..

OpenStudy (nincompoop):

LAUGHING MY ARSE OFF get out

OpenStudy (nincompoop):

d of u is 3(-3x+1)^2

OpenStudy (nincompoop):

or did I write it wrong?

OpenStudy (nincompoop):

and mic, I did it with "simplifying" and it is easier

OpenStudy (nincompoop):

it expanded it a little bit, but it is manageable and just follows one rule

OpenStudy (nincompoop):

try it

OpenStudy (anonymous):

so , for the f'' do I make de (u'v + uv') rule??

OpenStudy (anonymous):

uhh?

OpenStudy (anonymous):

so complicated

OpenStudy (shamil98):

Okay. i'm going to start over... f(x) = (1-3x)^3(2x-3) using the product rule, where u = (1-3x)^3 and v = 2x-3 you get \[f'(x) =(2x-3) \frac{ d }{ dx }[(1-3x)^3] + (1-3x)^3 \frac{ d }{ dx } (-3+2x)\]

OpenStudy (shamil98):

\[\frac{ d }{ dx }( u~v) = v\frac{ du }{ dx } + u \frac{ dv }{ dx }\] Now using the chain rule.. \[\frac{ d }{ dx } (1-3x)^3 = \frac{ du^3 }{ du } \frac{ du }{ dx }\] where u = 1-3x, and d/du (u^3) = 3u^2 \[f'(x) = (1-3x)^3 (\frac{ d }{ dx } (-3 + 2x)) + (-3+2x) ~~3(1-3x)^2\frac{ d }{ dx } (1-3x)\]

OpenStudy (anonymous):

okay thanks :):)

OpenStudy (nincompoop):

damn... laughing out loud

OpenStudy (shamil98):

nin wot

OpenStudy (nincompoop):

I think she already got the answer

OpenStudy (shamil98):

LOL

OpenStudy (anonymous):

No I'm just tired to do this fuging number haha :p..

OpenStudy (anonymous):

Finally got it ! hahaha

OpenStudy (e.mccormick):

Yes, that way worls well. There are a few you can use to get there, but that is a good choice.